### v3

# EEG-MACS: Manifold Attention and Confidence Stratification for EEG-based Cross-Center Brain Disease Diagnosis under Unreliable Annotations

Abstract.
Cross-center data heterogeneity and annotation unreliability significantly challenge the intelligent diagnosis of diseases using brain signals. A notable example is the EEG-based diagnosis of neurodegenerative diseases, which features subtler abnormal neural dynamics typically observed in small-group settings. To advance this area, in this work, we introduce a transferable framework employing Manifold Attention and Confidence Stratification (MACS) to diagnose neurodegenerative disorders based on EEG signals sourced from four centers with unreliable annotations. The MACS framework’s effectiveness stems from these features: 1) The Augmentor generates various EEG-represented brain variants to enrich the data space; 2) The Switcher enhances the feature space for trusted samples and reduces overfitting on incorrectly labeled samples; 3) The Encoder uses the Riemannian manifold and Euclidean metrics to capture spatiotemporal variations and dynamic synchronization in EEG; 4) The Projector, equipped with dual heads, monitors consistency across multiple brain variants and ensures diagnostic accuracy; 5) The Stratifier adaptively stratifies learned samples by confidence levels throughout the training process; 6) Forward and backpropagation in MACS are constrained by confidence stratification to stabilize the learning system amid unreliable annotations. Our subject-independent experiments, conducted on both neurocognitive and movement disorders using cross-center corpora, have demonstrated superior performance compared to existing related algorithms. This work not only improves EEG-based diagnostics for cross-center and small-setting brain diseases but also offers insights into extending MACS techniques to other data analyses, tackling data heterogeneity and annotation unreliability in multimedia and multimodal content understanding. We have released our code here: https://github.com/ICI-BCI/EEG-MACS.
EEG Signals, Cross-center Learning, Unreliable Annotation, Weakly-supervised Learning, Neurodegenerative Disease
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia ($M$ ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681645isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Modeling methodologiesccs: Applied computing Health care information systems
Figure 1. The proposed MACS framework tackles data heterogeneity and annotation unreliability in EEG-based brain disease diagnosis, aiming for superior performance across centers and validation in both neurocognitive and movement disorders.
The proposed MACS framework tackles data heterogeneity and annotation unreliability in EEG-based brain disease diagnosis, aiming for superior performance across centers and validation in both neurocognitive and movement disorders.
1. Introduction
Addressing the challenges posed by data heterogeneity and annotation unreliability is essential for the analysis and interpretation of multimedia and multimodal data (Wang et al., 2023b; Zhang et al., 2023b). These issues are particularly prevalent in human-centered computing, where they encounter even greater difficulties in small-group settings. A prime example is the analysis of neural dynamics, such as the automated diagnosis of cognitive and movement brain disorders through electroencephalogram (EEG) signals  (Wang et al., 2023a; Chang et al., 2023). While EEG-based diagnosis models have undergone extensive research, they are often disease-specific, rely on single-center studies, and depend on fully supervised learning with annotated data  (Yasin et al., 2021). Developing effective and transferable models capable of handling cross-center data and low-quality annotations is crucial for enhancing adaptability across different diseases and advancing EEG modeling techniques.
Current EEG analyzing methods range from traditional feature engineering to deep learning approaches. Various features extracted from domains like time (e.g., Hjorth parameters), frequency (e.g., Fourier Transform), dynamics (e.g., entropy), and functional networks (e.g., synchronization and phase coupling) have demonstrated efficacy due to EEG’s intrinsic characteristics (Coelho et al., 2023; Cavanagh et al., 2018; Suuronen et al., 2023; Heinrichs et al., 2023). The sparsity of handcrafted features has led to the adoption of deep learning models, such as convolutional neural networks, graph neural networks, and hybrid approaches, to capture spatial-temporal patterns, structured characteristics, and other high-level representations (Ju and Guan, 2023; Pan et al., 2022; Cai et al., 2023; Zhang et al., 2023a). However, these existing methods often struggle with noisy data and lack cross-center applicability. This may primarily be due to: 1) the ineffective mapping that cannot project EEG signals into a high-level feature space invariant across diverse data distributions from multiple centers; 2) the absence of integrating advanced learning strategies to tackle complex learning issues under unreliable annotations effectively.
From this perspective, our work presents a novel framework aimed at developing a transferable model adept at managing unreliable annotations, as illustrated in Figure 1. This model specifically targets the recognition of neurodegenerative disorders. These include neurocognitive disorders, represented by mild cognitive impairment(MCI) and Alzheimer’s Disease(AD), and movement disorders, exemplified by Parkinson’s Disease(PD). We selected these three diseases due to the potential progression relationship between AD and MCI (Knopman et al., 2021) as well as the observed comorbidity between PD and cognitive impairment (Dag et al., 2021; Anjum et al., 2024). The proposed framework is characterized by its innovative use of Manifold Attention and Confidence Stratification (MACS), which optimally synergizes modules including the Augmentor, Switcher, Encoder, Projector, and Stratifier. MACS’s success in achieving transferability across centers and diseases under unreliable annotations can be attributed to three key factors: 1) The establishment of an optimized mapping that leverages the strengths of both Euclidean and Riemannian manifold spaces enables the model to extract more effective EEG representations. 2) The integration of supervised and self-supervised learning strategies through confidence stratification to address the issue of learning from unreliable annotations. 3) An effective encoder enhances the representation space, which improves confidence stratification accuracy, thereby enabling the encoder to develop more robust and inductive representations in a mutually reinforcing cycle. Our primary contributions are outlined below:
  • Introducing a novel EEG-based framework, MACS, designed for learning from unreliable annotated EEG signals with cross-center transferability.
    Augmentor::
    Enriches the data space via augmentation.
    Switcher::
    Mitigates overfitting on incorrect labels.
    Encoder::
    Integrates Manifold and Euclidean spaces to improve EEG representation learning.
    Projector::
    Features dual heads to assess both data consistency and diagnostic accuracy.
    Stratifier::
    Stratifies data based on confidence levels.
  • Establishing confidence stratification-based constraints for MACS’s forward and back-propagation that create a self-organizing system to enhance representation learning and confidence assessment, thereby fostering a virtuous cycle.
  • Demonstrating MACS’s superior performance in learning from unreliable annotations compared to state-of-the-art (SOTA) methods, validated across two types of diseases and through cross-center testing and fine-tuning.
  • Making the code public to contribute multimedia community.
Figure 2. Overview of the MACS framework and subject-independent experiments across diseases and centers. The Projector has a dual-head structure for latent space representation and classification. The Stratifier categorizes samples by label confidence, constraining brain variants for contrastive learning, and discriminative loss for distrusted samples. Refer to Figures 3 and 4 for Switcher and Encoder, respectively.
2. Related Work
2.1. Learning with Unreliable Annotations
Noisy label learning contains model-free and model-based strategies for reducing the influence of incorrect labels. These strategies involve estimating noise patterns and conducting supervised learning with clean samples (Hendrycks et al., 2018), as well as handling noisy labels and refining the model by addressing internal conflicts (Malach and Shalev-Shwartz, 2017; Tan et al., 2021). Such methods depend on the analysis of the relationships within noisy data. Recent research in time-series (Ma et al., 2023) and image domains (Xiao et al., 2023; Li et al., 2022) applies unique principles to delineate relationships. In the image domain, Promix (Xiao et al., 2023) adopts small-loss criteria and prediction consistency to filter high-confidence examples, followed by learning through Debiased Semi-Supervised Training. Sel-CL (Li et al., 2022) employs nearest neighbors to select confident pairs for supervised contrastive learning. In the time-series domain, CTW (Ma et al., 2023) selects confident examples based on small-loss criteria and applies time-warping to these instances to learn more robust representations. We use these to benchmark MACS and, following (Li et al., 2020; Jiang et al., 2020), incorporate mix-up techniques (Berthelot et al., 2019; Zhang et al., 2021) to enhance learning from noisy labels.
2.2. Contrastive Learning for Time Series
Contrastive learning has demonstrated effectiveness in both unsupervised (e.g., SimCLR (Chen et al., 2020)) and supervised (e.g., SupCon (Khosla et al., 2020)) contexts. It provides valuable self-learning strategies for pre-training in time series data, such as enhancing time-frequency (Zhang et al., 2022a) and cross-sample temporal consistency (Wang et al., 2023a). However, the impact of such pre-training on handling unreliable annotated data is less pronounced. Augmentation-based contrast (Eldele et al., 2023) shows promise for semi-supervised learning with missing labels but is less effective without true label priors. Therefore, our proposed MACS introduces multi-view contrastive learning to overcome these limitations, guided by estimated confidence levels.
2.3. Manifold Learning for EEG Signals
Manifold-based modeling excels in EEG signal analysis for brain-computer interfaces (BCI) (Yger et al., 2016; Chevallier et al., 2021), leveraging Riemannian geometry for high-dimensional neural data representation through affine-invariant geometric distances (Huang and Van Gool, 2017; Sabbagh et al., 2019). Using EEG’s Riemannian structure for BCI domain-adaptation (Ju and Guan, 2023; Kobler et al., 2022; Bonet et al., 2023) and multi-task application (Pan et al., 2022) indicates the potential of Manifold geometry learning in identifying robust latent spaces. Such evidence bolsters the Encoder’s design in MACS. Our work advances current geometry learning by integrating dynamic functional networks with cross-attention mechanisms. Additionally, findings from our pilot study suggest that when integrated with contrastive learning, Riemannian geometry learning offers a promising alternative to GNNs (Zhang et al., 2022b).
3. MACS
In this section, we elucidate the methodology for learning from brain signals annotated with unreliable labels using the Manifold Attention and Confidence Stratification (MACS) framework. As depicted in Figure 2, MACS handles unreliable labels and enhances cross-center transferability through its modules (in Sec. 3.1), adaptive constraints (in Sec. 3.2), and training objective(in Sec. 3.3).
Assuming that brain activity is monitored using $d$ sensors at a sampling frequency of $f_{s}$ for a duration of $t$ seconds, we obtain the following set of observations as the input for the MACS framework:
(1) $\bm{A}_{d,T}^{(n)}=\begin{pmatrix}a_{1,1}&a_{1,2}&\cdots&a_{1,T}\\ a_{2,1}&a_{2,2}&\cdots&a_{2,T}\\ \vdots&\vdots&\ddots&\vdots\\ a_{d,1}&a_{d,2}&\cdots&a_{d,T}\end{pmatrix}$,
where $T=f_{s}\cdot t$ denotes the total number of sampling points, and $n\in\{1,2,\ldots,N\}$ represents a specific individual from a total of $N$ subjects under observation.
The MACS framework identifies brain states as $\hat{Y}^{(n)}=F_{\Theta}(\bm{A}_{d,T}^{(n)})$ in scenarios with unreliable annotations ${Y}^{(n)}$, utilizing five key components: 1) Augmentor, 2) Switcher, 3) Encoder, 4) Projector, and 5) Stratifier. The initial four modules are integral to the framework’s forward propagation, while the Stratifier systematically regulates the learning mechanism.
3.1. MACS Modules
Augmentor: Producing Brain Variants through Dual Random Transform.
In this module, data augmentation is applied to the preprocessed brain data as described in (Zhang et al., 2022a), following the equation:
(2) $f_{Ag}(A_{d,T_{s}},\sigma)=A_{d,T_{s}}+\bm{\epsilon},$
where $\bm{\epsilon}_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is independently and identically distributed, with $\sigma$ representing the standard deviation. Additionally, $A_{d,T_{s}}$ is a segmented, non-overlapping fragment of $\bm{A}_{d,T}$, utilized to reduce computational load and increase the system’s robustness. Two random transformations are concurrently executed, resulting in dual brain variants. Differing from the method of specific weak and strong augmentations for two views (Eldele et al., 2021), our approach operates a random configuration, leading to improved performance.
Figure 3. The Switcher employs conditional interpolated blending for trusted samples and bypasses distrusted ones to mitigate overfitting on incorrectly labeled samples.
Switcher: Blending-based Variant Generation Conditional on Confidence Levels.
In the Switcher module, dual brain variants undergo selective processing. As in Figure 3, distrusted inputs are directly forwarded to subsequent modules, while trusted ones are routed to a blender for generating interpolated samples. Inspired by the benefits of sample and network mixing, as discussed in (Berthelot et al., 2019) and (Li et al., 2020), and aiming to enhance representation learning with partial labels and reduce confirmation bias in noisy data, we distort each trusted sample $A_{d,T_{s}}^{*}$ and its corresponding label $y^{*}$ by interpolating it with another randomly selected sample $(A_{d,T_{s}}^{+},y^{+})$, to produce blended variants, as described in Eq.(3). $\lambda$ is sampled from a uniform Beta(1, 1) distribution and adjusted to ensure it is always at least 0.5 by setting it to the greater value between $\lambda$ and $1-\lambda$.
(3) $f_{Sw}((A_{d,T_{s}}^{*},y^{*}),\lambda)=\lambda(A_{d,T_{s}}^{*},y^{*})+(1-\lambda)(A_{d,T_{s}}^{+},y^{+})$
This choice is implemented to balance the retention of the primary input’s information and the introduction of diversity from the randomly indexed input throughout the mixing process.
Differing from the referenced studies (Berthelot et al., 2019; Li et al., 2020), our method presents two distinctions: 1) We produce dual-blended brain variants while employing shared network modules, thus effectively decreasing model complexity. 2) We initiate data blending exclusively on samples exhibiting higher confidence levels, aiming to mitigate concerns related to unreliable labels. The confidence level associated with each sample in the mini-batch is evaluated using the Stratifier module in the MACS framework.
Encoder: Mapping EEG Dynamics onto Riemannian Manifold for Attention-Based Analysis.
The Encoder module $f_{En}$, grounded in our feature engineering findings (Section 4.3), initiates with a convolution starter $g_{str}$. This component reduces noise/artifacts and captures specific wave patterns (e.g., theta, alpha rhythms) and time-domain features of brain activity using convolutions and normalization, effectively decoding brain signals at a lower level(Défossez et al., 2023). This process yields $X_{d,T_{s}}=g_{\text{str}}(A_{d,T_{s})}$. Subsequently, a temporal clipper $g_{\text{clp}}$ is applied to these low-level features $X_{d,T_{s}}$ across the sampling dimension. This operation enhances their representation by leveraging the high temporal resolution of EEG signals.A series of feature clips $\{\ X_{dim_{1},t_{i}}\}\ _{i=1}^{I}$ are then processed by a synchronization analyzer $g_{syn}$, which measures the coupling between pairs of $d$ sensors, ultimately producing a dynamic functional network represented as $\Phi=\left[\phi_{t_{1}},\ldots,\phi_{t_{i}},\ldots,\phi_{t_{I}}\right]$.
Figure 4. The Encoder combines Riemannian and Euclidean metrics for feature extraction, leading to a manifold-based attention mechanism that effectively captures characteristics of spatiotemporal complexity and dynamic synchronization.
With evidence from learning Riemannian geometry for structured brain data as shown in (Bonet et al., 2023; Pan et al., 2022), a manifold converter $g_{mfd}$ transforms dynamic network $\Phi$ into a set of symmetric positive definite (SPD) matrices, $S_{d}^{++}(\mathbb{R})$, thus defining a Riemannian manifold $\mathcal{M}$. This transformation involves eigenvalue decomposition and transpose manipulation of each adjacency matrix, represented as $\mathcal{D}=g_{mfd}(\Phi)$. To measure those $S_{d}^{++}$, we utilized the Log-Euclidean Riemannian metric (Huang and Van Gool, 2017), as it serves as the first-order approximation of the Affine-Invariant metric. At each point $\mathcal{D}_{t_{i}}$ on the manifold $\mathcal{M}$, a tangent space is defined by $\log:\mathcal{M}\to\mathcal{T}_{\mathcal{M}}$, where the inverse is operated by exponential mapping, as illustrated in Figure 4. The distance between any two points on the manifold $\mathcal{M}$ is then calculated using the following equation:
(4) $d(\mathcal{D}_{t_{i}},\mathcal{D}_{t_{j}})=\left\lVert\log(\mathcal{D}_{t_{i}})-\log(\mathcal{D}_{t_{j}})\right\rVert_{F}^{2}.$
Utilizing this metric to reflect the correlation between points $\mathcal{D}_{t_{i}}\in S_{d}^{++}$, we constructed a manifold-based dynamic attention block $g_{datt}$, following the approach in (Pan et al., 2022), but with modifications to incorporate a cross-temporal attention mechanism. This integration of dynamic relationships is mathematically represented as:
(5) $\mathcal{F}_{i}=\sum_{j=1,j\neq i}^{I}\delta\left(d(f_{W_{k}}(\mathcal{D}_{t_{i}}),f_{W_{q}}(\mathcal{D}_{t_{j}}))\right)\cdot\log(f_{W_{v}}(\mathcal{D}_{t_{i}})),$
where $\delta$ scales the values to a range of 0 to 1. $f_{W_{k}}$ represents the operation of $W_{k}$ multiplied by a matrix and then by $W_{k}^{\top}$. This can be expressed as $f_{W_{k}}(X)=W_{k}XW_{k}^{\top}$. Similarly, $f_{W_{q}}$ and $f_{W_{v}}$ operate in the same manner, where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d_{1}}$ are learned weight matrices used for the bilinear mapping. Subsequently, the fused feature matrices $\mathcal{F}=\{\mathcal{F}_{i}\}$ are subjected to a nonlinear activation process, involving the rectification of eigenvalues, as detailed in (Huang and Van Gool, 2017). To enable the measurement of these features $\mathcal{F}$ in Euclidean space for further processing, the embeddings $\hat{\mathcal{F}}$ are obtained by mapping $\mathcal{F}$ onto the tangent space, followed by flattening and concatenating them together.
Projector: Dual-Branch Mapping for Latent Distribution and Discriminative Classification.
[…]
Stratifier: Measuring Latent Similarity for Confidence Stratification.
This module $f_{St}$ plays a crucial role in assessing the reliability of the learned representations, which in turn influences the learning mechanism of the MACS framework.
It categorizes the confidence levels of samples based on $\mathcal{Z}$ through the following steps:
  • Measuring the cosine similarity between samples in $\mathcal{Z}$.
    (6) $c\left(z_{i},z_{j}\right)=\frac{z_{i}z_{j}^{\top}}{\left\|z_{i}\right\|\left\|z_{j}\right\|}$
  • Locating the K closet neighbors of each sample $z_{i}$ in $\mathcal{Z}$ using the k-Nearest Neighbors method.
  • Obtaining a neighbor-determined label $\bar{y}$ for each sample $z_{i}$ by averaging the original labels of its neighbors.
    (7) $p_{c}\left(z_{i}\right)=$ $\frac{1}{K}\sum_{\begin{subarray}{c}k=1\\ z_{k}\in\mathcal{N}_{i}\end{subarray}}^{K}\mathds{1}_{y_{k}=c},c\in\{0,1\},$
    $\bar{y}=$ $\arg\max_{c}p_{c}(z_{i}),$
    where $\mathcal{N}_{i}$ denotes the neighbourhood of K samples to $z_{i}$. The indicator function, denoted as $\mathds{1}_{B}$, is a function that returns $1$ when the condition $B$ is satisfied and 0 otherwise.
  • Considering a sample trusted if its neighbor-determined label matches its original label. Following (Ortego et al., 2021), we also adopted a dynamic thresholding approach to ensure class balance.
  • Identifying trusted pairs $\Psi$ from trusted samples $\mathcal{Z}_{\text{crd}}$ based on label identity.
For the distrusted samples $\mathcal{Z}_{\text{dst}}=\mathcal{Z}\setminus\mathcal{Z}_{\text{crd}}$, corresponding auxiliary labels are generated through the forward-propagation process $\bm{A}\xrightarrow{f_{En}}\hat{\mathcal{F}}\xrightarrow{f_{Pc}}\hat{Y}$.
Algorithm 1 MACS
[…]
2:Learned Model $\Theta$
[…]
4:for $t=1,2,\ldots,T_{\text{max}}$ do
5:  Stratifier: Identify trusted examples $A_{\text{tru}}$ and distrusted examples $A_{\text{dst}}$ based on the consistency between $\mathcal{Z}=f_{Pl}(f_{En}(A_{d,T_{s}}))$ and provided unreliable labels $Y$
6:  Augmentor: $f_{Ag}(A_{d,T_{s}},\sigma)=A_{d,T_{s}}+\bm{\epsilon}$, generate two brain variants
7:  if $\text{samples}\in A_{\text{dst}}$ then
8:   Encoder: $\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
[…]
10:   Loss: Compute self-supervised contrastive loss $\mathcal{L}^{Ag}$ and use softmax prediction $\hat{Y}$ without data augmentation to compute discriminative loss $\mathcal{L}^{DL}$
11:  else
12:   Switcher: $f_{Sw}((A_{d,T_{s}}^{*},y^{*}),\lambda)=\lambda(A_{d,T_{s}}^{*},y^{*})+(1-\lambda)(A_{d,T_{s}}^{+},y^{+})$
13:   Encoder: $\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
[…]
15:   Loss: Compute mixed self-supervised loss $\mathcal{L}^{Sw}$, supervised contrastive loss $\mathcal{L}^{St}$, and use original label $Y$ without data augmentation to compute discriminative loss $\mathcal{L}^{DL}$
16:  end if
17:end for
3.2. MACS Constraints
MACS constraints are dependent on the confidence stratification adaptively assessed by the Stratifier module during training, which is pivotal in stabilizing representation learning in the context of unreliably annotated data. Intuitively, they regulate the behavior of the modules, influencing the forward propagation of features and the computation of backward losses, as depicted in Figure 2. Four types of constraints are implemented within the framework: they are applied to both the Augmentor and Switcher modules and also influence the multi-view contrastive loss and discriminative loss. Those constraints work as follows:
1) The first constraint specifically controls the gradients of Switcher module $\nabla f_{Sw}$. It effectively utilizes linearly interpolated samples to enrich data associated with trusted labels while simultaneously mitigating the risk of being misled by unreliable data.
2) The second constraint type facilitates a ’divide and conquer’ approach within the multi-view contrastive loss. Specifically, contrastive learning is initiated within three groupings: among trusted pairs $\Psi$, among blended variants of trusted samples $\mathcal{Z}_{\text{crd}}$, and among dual random brain variants $\mathcal{Z}$.
3) The third constraint directs the Augmentor module $f_{Ag}$ to bypass distrusted samples and put them into the Encoder and Projector to generate auxiliary labels, providing alternatives to unreliable labels for backpropagation.
4) The fourth constraint conditionally influences the discriminative loss. It evaluates the errors of trusted samples using their original labels, while the gradients for distrusted samples are computed based on auxiliary labels.
3.3. MACS Training Objective
The overall training objective, as formulated in Eq.(8), is an equal sum of contrastive and discriminative losses.
(8) $\mathcal{L}=$ $\mathcal{L}^{CL}+\mathcal{L}^{DL}$
3.3.1. Multi-view Contrastive Loss
In our framework, the contrastive loss $\mathcal{L}^{CL}$ consists of three components: $\mathcal{L}^{Ag}$, $\mathcal{L}^{Sw}$, and $\mathcal{L}^{St}$, each addressing different aspects of the data. All components adhere to the core formulas defined in Eq.(9). The computation involves aggregating $C_{i,j}$ for all pairs of samples $x_{j}$ within the same minibatch that share the same label as $x_{i}$ (i.e., $y_{i}=y_{j}$), across all $N_{y_{i}}$ such samples. Each minibatch contains $N_{m}$ samples, excluding self-contrast cases where $i=j$. A temperature parameter $\tau$ modulates the scale of the dot products in the softmax denominator.
(9) $\mathcal{L}^{\mathcal{O}}_{i}(z_{i},y_{i})=$ $-\frac{1}{2N_{y_{i}}-1}\sum_{j=1}^{2N_{m}}\mathds{1}_{i\neq j}\mathds{1}_{y_{i}=y_{j}}\mathcal{C}_{ij},$
$\mathcal{C}_{ij}=$ $\log\left(\frac{\exp(z_{i}\cdot z_{j}/\tau)}{\sum_{r\in\{1,...,2N_{m}\}\setminus\{i\}}\exp(z_{i}\cdot z_{r}/\tau)}\right).$
1) $\mathcal{L}^{Ag}$: It aims to minimize the distance between dual augmentations for distrusted examples $\mathcal{Z}_{\text{dst}}$ , which constitutes a self-supervised learning approach.
2) $\mathcal{L}^{Sw}$: This targets the enhancement of similarity among blending variants, but exclusively for trusted samples $\mathcal{Z}_{\text{crd}}$, which constitutes a mixed self-supervised learning approach.
3) $\mathcal{L}^{St}$: This component incorporates interpolation supervised contrastive learning techniques, as exploited in (Ortego et al., 2021; Li et al., 2022), and shares the same hyperparameter $\lambda$ with Eq.(3). It effectively combines $\mathcal{L}^{St}_{i\in\Psi^{*}}$ and $\mathcal{L}^{St}_{i\in\Psi^{+}}$ from trusted pairs $\Psi=\Psi^{*}\cup\Psi^{+}$. It can be expressed as $\mathcal{L}_{i}^{St}=\lambda\mathcal{L}_{i}\left(z_{i},y^{*}\right)+(1-\lambda)\mathcal{L}_{i}\left(z_{i},y^{+}\right)$.
3.3.2. Discriminative Loss
This loss function plays a direct role in quantifying the recognition of abnormalities. It is conditionally structured as detailed in Eq.(10). In this equation, $Y^{(n)}$ represents the original label, $\hat{Y}^{(n)}$ denotes the generated auxiliary label. Algorithm 1 presents the pseudocode for MACS.
(10) $\mathcal{L}^{DL}_{i}=$ $-\lambda\tilde{Y}^{(*)T}\log(\mathcal{E}^{(i)})$
$-(1-\lambda)\tilde{Y}^{(+)T}\log(f_{Pc}(\mathcal{E}^{(i)}),$
[…]
[MCI] 4-Fold Cross-Validation [PD] 3-Fold Cross-Validation Comprehensive Average
Feature Domains Accuracy F1 Accuracy F1 Accuracy F1
Frequency(Cavanagh et al., 2018) 58.64(2.06) 66.03(1.47) 74.07(0.07) 71.90(0.02) 66.36 71.51
Entropy(Suuronen et al., 2023) 63.07(1.46) 69.89(1.50) 61.11(0.21) 57.68(2.29) 62.09 72.37
Hjorth(Coelho et al., 2023) 67.05(0.80) 68.37(0.75) 64.81(0.27) 63.78(0.49) 65.93 64.65
Functional Connectivity-Corr(Lopez et al., 2023) 65.34(1.64) 67.56(2.58) 66.67(0.21) 61.39(0.78) 66.01 63.99
Functional Connectivity-SPD (Bonet et al., 2023) 64.43(0.77) 68.18(0.99) 79.63(0.27) 79.46(0.17) 72.03 74.80
Table 1. Feature engineering assesses the effectiveness of various EEG domains in neurodegeneration recognition.
Table 2. Comparative study compares the SOTA methods for learning from both well-annotated and unreliable data.
[MCI] 4-Fold [PD] 3-fold
Scenarios Methods Accuracy F1 Accuracy F1 Overall Accuracy
(Ju and Guan, 2023) [Ju et al., TNNLS, 2022] Tensor-CSPNet 80.78(7.93) 77.70(11.63) 75.92(6.41) 73.20(8.16) 78.35(3.44)
(Pan et al., 2022) [Pan et al., NeurIPS, 2022] MAtt 81.97(3.96) 79.56(8.29) 79.63(11.56) 83.07(7.56) 80.80(1.65)
(Wang et al., 2023a) [Wang et al., NeurIPS, 2023] COMET 73.25(26.02) 70.09(24.82) 75.47(9.75) 74.92(14.17) 74.36(1.57)
Clean Annotation Encoder in MACS Final Epoch 87.65(4.31) 86.47(7.73) 85.18(8.48) 83.46(11.67) 86.42(1.75)
(Ma et al., 2023)[Ma et al., IJCAI, 2023] CTW 75.40(10.28) 76.75(9.80) 75.93(3.21) 78.73(2.20) 75.67(0.37)
(Xiao et al., 2023)[Xiao et al., IJCAI, 2023] Promix 55.14(18.94) 56.17(11.49) 79.63(8.49) 77.34(10.97) 67.39(17.32)
(Ma et al., 2023) [Ma et al., IJCAI, 2023]
(Xiao et al., 2023) [Xiao et al., IJCAI, 2023]
CTW Encoder + Promix 79.75(4.69) 80.83(7.21) 81.48(6.41) 80.25(8.15) 80.62(1.22)
(Li et al., 2022) [Li et al., CVPR, 2022] Sel-CL 63.29(22.32) 54.82(19.42) 66.67(5.56) 65.42(4.49) 64.98(2.39)
(Li et al., 2022) [Li et al., CVPR, 2022] Sel-CL+ 74.16(7.73) 73.41(11.98) 57.41(21.03) 43.06(37.35) 65.79(11.84)
(Xiao et al., 2023)[Xiao et al., IJCAI, 2023] Encoder for Promix 85.38(4.41) 85.56(4.90) 83.33(5.56) 81.05(8.49) 84.36(1.45)
(Li et al., 2022) [Li et al., CVPR, 2022] Encoder for Sel-CL 72.08(12.76) 70.05(9.37) 83.33(5.56) 81.51(9.04) 77.71(7.95)
(Li et al., 2022) [Li et al., CVPR, 2022] Encoder for Sel-CL+ 85.43(4.16) 84.18(7.01) 83.33(5.56) 81.05(8.49) 84.38(1.48)
Unreliable Annotation MACS Final Epoch 88.74(4.61) 88.18(7.23) 87.04(3.21) 86.40(1.90) 87.89(1.20)
4. Experiments
We conduct four subject-independent experiments (presented in Figure 2) to validate MACS’s performance in learning from unreliable annotations and transferability: 1) Three-fold cross-validation using public PD data from Center A (Model $\Theta_{PD}$); 2) Four-fold cross-validation with MCI (MCI due to AD) data collected from Center C (Model $\Theta_{MCI}$); 3) Direct testing $\Theta_{PD}$ on a cross-center PD corpus (Center B) for transferability evaluation; 4) Fine-tuning $\Theta_{MCI}$ on a cross-center AD corpus (Center D) for transferability assessment. Please refer to Appendix for implementation details.
4.1. Experimental Datasets
PD Datasets
The PD data includes two sets. The first set, from the University of New Mexico(NMU), comprises data from 27 patients and 27 controls (Cavanagh et al., 2018). The second set, from the University of Iowa(IU), includes data from 14 patients and 14 controls (Anjum et al., 2020). Both datasets were acquired using 64-channel Ag/AgCl electrodes with the Brain Vision system at a sampling rate of 500 Hz.
MCI Dataset
The MCI dataset, from a local hospital in City C, includes data from 46 MCI patients and 43 age-matched normal controls (NC), recorded in an eye-closed state. Data was captured using a 64-channel Ag/AgCl electrodes Brain Product system at a sampling rate of 5000 Hz.
AD Dataset
The AD dataset, from a local hospital in City D, comprises data from 20 AD patients and 20 age-matched NC volunteers. Each participant provided two EEG samples (eye-closed and eye-open states), recorded using a 16-channel Ag/AgCl electrodes Symtop system at 1024 Hz.
4.2. Training Configuration
For each dataset, we created a set of unreliable labels, assigning incorrect labels to a percentage of the samples defined by $\alpha$. Performance was evaluated using true labels, verified, and corrected by experts. Results for $\alpha=0.3$ are presented as main findings, while those for $\alpha=0.5$ are detailed in the Appendix. The MACS model was trained over 30 epochs using Stochastic Gradient Descent with a momentum of 0.9 and a weight decay of $10^{-4}$. A step-learning rate scheduler improved training efficiency by starting at 0.1 and reducing it by a factor of 0.1 every 10 epochs. Similarly, a three-fold cross-validation was executed on the PD dataset from the NMU center, with a batch size of 60 We conducted four-fold cross-validation on the MCI dataset with a batch size of 128, and three-fold cross-validation on the PD dataset from the NMU center with a batch size of 60. The MACS framework was developed on a Linux-based system using PyTorch (version 2.0.0), with hardware including a GeForce RTX 3090 GPU and an Intel i9-12900K CPU.
4.3. Feature Engineering Study
Feature engineering insights have led to the development of the MACS Encoder. EEG frequency markers (Cavanagh et al., 2018)are inductive for PD but found less effective for MCI. EEG complexity measured by the entropy(Suuronen et al., 2023) and Hjorth parameter(Coelho et al., 2023) enhanced MCI detection but was less effective than frequency domain features in PD cases. We compared the functional connectivity characterized by Pearson correlation (Corr) in Euclidean space (Lopez et al., 2023) and SPD in manifold space (Bonet et al., 2023) and found the advantage of manifold representations. This evidence has informed the integration of a specialized block in the Encoder to process underlying temporal-spectral-spatial features and analyze synchronization characteristics through manifold geometry.
4.4. Comparison Study
We benchmarked MACS against SOTA methods, evaluating its performance in scenarios with unreliable annotations and independently assessing the Encoder in scenarios with clean annotations (see Table 2). For the clean data learning comparison, we employed advanced EEG encoders that target manifold geometry(Ju and Guan, 2023), integrate attention mechanisms(Pan et al., 2022), and utilize contrastive learning(Wang et al., 2023a). Given the absence of a dedicated SOTA model for EEG data with unreliable annotations, we compared MACS against leading methods in the time series(Ma et al., 2023) and natural image domain(Xiao et al., 2023; Li et al., 2022). Specifically, we replaced their encoders with our Encoder in MACS to validate the efficacy of our combined manifold-Euclidean representation learning strategy for EEG signals. Overall, MACS’s weakly supervised learning mechanism, combined with an effective Encoder, enables superior comprehensive performance across two disease types. We unveil MACS’s learning process through t-SNE mapping in Figure 5 (see Appendix for the PD case). Observations indicate that MACS gradually selects the correct trust samples, enhancing representation learning and forming class-specific clusters.
4.5. Hyperparameter Tuning Study
Memory Length
The multi-view contrastive loss is crucial for the MACS framework. Effective contrastive learning heavily depends on the adequacy of positive and negative pairs; therefore, inspired by (Ortego et al., 2021), we proposed maintaining a memory of previous steps to circumvent the constraints imposed by batch size, enabling more extensive use of available small-scale data. For example, we summarize the research findings for MCI in Table 3, suggesting that memory length contributes to learning efficiency, but its relationship is not monotonic. Please note that the methods involving contrastive learning are discussed in Table 2, where we also determined their optimal performance to ensure a fair comparison.
Table 3. Investigation of hyper-parameters in MACS.
Parameters Configurations Accuracy F1
0 83.10(6.93) 83.64(9.09)
200 87.60(4.46) 86.39(7.58)
300 88.74(4.61) 88.18(7.23)
400 86.31(5.55) 84.71(9.69)
Memory Length 500 85.33(4.52) 84.07(7.77)
2s 88.74(4.61) 88.18(7.23)
1s 87.65(4.31) 86.75(7.87)
Temporal Scale 500ms 74.06(6.09) 77.60(8.15)
Temporal Scale Temporal scale refers to the time interval for constructing dynamic functional networks in the Encoder. The findings indicate that a second-level temporal scale is crucial for capturing patterns. For example, results from the MCI dataset are included in Table 3. More effective EEG markers that appear at larger temporal scales may be attributed to the ’slowing’ phenomenon in brain activity, which is more prominently manifested in patients with neurodegenerative disorders (Song et al., 2018; Jones et al., 2022).
Table 4. Ablation study of the components in MACS.
[MCI] 4-Fold [PD] 3-Fold
Configuration Accuracy F1 Accuracy F1
MACS 88.74(4.61) 88.18(7.23) 87.04(3.21) 86.40(1.90)
w/o Memory 83.10(6.93) 83.64(9.09) 85.18(3.21) 85.15(3.37)
w/o Augmentor 84.24(5.95) 82.17(11.57) 85.18(3.21) 85.32(1.89)
w/o Switcher 87.55(6.89) 86.05(9.79) 83.33(5.56) 81.05(8.49)
w/o Confidence
84.19(9.62) 84.68(12.24) 79.63(6.41) 79.46(5.01)
w/o Auxiliary
81.92(7.58) 82.23(9.72) 83.33(0.0) 82.20(2.10)
w/o MA 84.39(7.31) 83.99(5.98) 79.63(3.20) 76.88(6.59)
w/o CS 70.95(10.53) 68.13(9.39) 77.78(9.62) 75.08(13.02)
Figure 5. Qualitative comparison of MACS with SOTA frameworks using t-SNE visualization based on latent distribution.
Figure 6. Evaluation of MACS’s Cross-center Transferability: (a) Direct testing results of MACS, trained on Center A’s PD data applied to Center B’s PD data; (b) Fine-tuning testing results of MACS, trained on Center C’s MCI data applied to Center D’s AD data, using only a limited percentage of labels.
4.6. Ablation Study
Table 4 outlines the impacts of MACS’s components, where ’w/o Memory’ denotes omitting large-scale storage in contrastive learning. The findings on Augmentor highlight the importance of comparing dual brain variants. The variants enhanced by interpolation in Switcher significantly benefit small-scale clinical data. The Stratifier’s necessity was confirmed by evaluating classification loss using all data in mini-batch without considering confidence levels (denoted as ’w/o Confidence’), and by giving discriminative feedback only for trusted samples, omitting auxiliary input for distrusted samples (denoted as’w/o Auxiliary’). Furthermore, we implemented cross-entropy loss in the model devoid of confidence stratification, marked as ’w/o CS’, to validate our hypothesis about this mechanism’s effectiveness.
4.7. Transferability Evaluation
From cross-validation tests on PD(NMU) and MCI datasets, we derived N-fold MACS models for PD ($\Theta_{PD}$) and MCI ($\Theta_{MCI}$). We assessed MACS’s transferability by testing $\{\theta_{PD}^{i}\}_{i=0}^{2}$ on a cross-center PD corpus (from IU), and fine-tuning $\{\theta_{MCI}^{i}\}_{i=0}^{3}$ on a cross-center AD corpus. To address the spatial resolution gap between MCI and AD data, we employed EEG’s interpolation technique (Gil Ávila et al., 2023). The averaged evaluation results from $\Theta_{PD}$, halted at both optimal and final epochs (Figure 6(a)), and successful fine-tuning on AD with few labels (Figure 6(b)), underscore MACS’s notable transferability in recognizing neurodegenerative disorders via EEG.
5. Conclusion
This work introduces MACS, a framework that leverages Manifold Attention and Confidence Stratification to address data heterogeneity and annotation unreliability in EEG modeling. Specifically, MACS fuses Euclidean space with manifold geometry to enhance representations and tailors forward and back-propagation based on stratified confidence levels. We have demonstrated superior performance in recognizing two types of diseases and further validated MACS’s transferability through cross-center testing and fine-tuning. The techniques used in MACS, effective in signal modeling under unreliable labels, could inspire advances in EEG-based cognitive and emotional computing, as well as in other areas of multimedia representation learning involving low-quality annotations.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No. 62306089), the Shenzhen Science and Technology Program (Grant No. RCBS20231211090800003), and the China Postdoctoral Science Foundation (Grant Nos. 2023M730873 and GZB20230960).
Appendix A Data Details
A.1. EEG Acqusition
Mild Cognitive Impairment (MCI) Dataset
The MCI dataset was acquired from a local hospital using a 64-channel Brain Products EEG system. This system was configured as Figure A1 with the following electrodes: Fp1, Fpz, Fp2, AF7, AF3, AFz, AF4, AF8, F7, F5, F3, F1, Fz, F2, F4, F6, F8, FT7, FC5, FC3, FC1, FCz, FC2, FC4, FC6, FT8, T7, C5, C3, C1, Cz, C2, C4, C6, T8, TP9, TP7, CP5, CP3, CP1, CPz, CP2, CP4, CP6, TP8, TP10, P7, P5, P3, P1, Pz, P2, P4, P6, P8, PO7, PO3, POz, PO4, PO8, O1, Oz, O2, and Iz. The electrodes at AFz and FCz were designated as ground and reference, respectively, and were excluded from the data analysis. A sampling rate of 5000 Hz was utilized during the EEG recording process. Data from the ground and reference channels were excluded from the dataset. A total of one hundred participants with cognitive decline were initially recruited. Following the inspection of medication history and other case-relevant factors (conducted by clinical physicians) as well as strict age-matching with the control group (carried out by engineering personnel), a final cohort comprising 46 subjects in the MCI group and 43 subjects in the HC group was deemed eligible for the study. Participants were instructed to sit comfortably and maintain a state of eyes-closed relaxation for an 8-minute recording session.
Alzheimer’s Disease (AD) Dataset
The AD dataset was procured from clinical monitoring at an external center, utilizing a 16-channel Symtop EEG amplifier. This system employed the International 10-20 system for electrode distribution, a configuration derived from the more comprehensive 10-10 system but with fewer channels, specifically: Fp1, Fp2, F7, F3, Fz, F4, F8, T7, C3, Cz, C4, T8, P7, P3, Pz, P4, P8, O1, and O2. The electrodes utilized for the AD data collection are marked in a darker color in Figure A1. Data capture for each participant lasted one minute, operating at a sampling frequency of 1,024 Hz. The linked earlobes, A1 and A2, served as the reference points. Under expert supervision, each participant contributed two samples—one with eyes closed and the other with eyes open. Each sample was ensured to have a minimum effective duration of 8 seconds, devoid of any significant artifacts or poor recording intervals.
Figure A1. Topographic mapping of electrode positions utilized for EEG dataset collection. The electrodes F9, F10, FT9, FT10, P9, P10, PO9, PO10, O9, and O10 were not included in the 64-channel Brain Products EEG system used for the Mild Cognitive Impairment (MCI) dataset collection. Additionally, the electrodes Fpz, F9, F10, P9, P10, PO9, PO10, O9, O10, and Iz were excluded from the 64-channel Brain Vision EEG system used for the collection of two Parkinson’s Disease (PD) datasets. The electrodes marked in darker color indicate inclusion in the 16-channel Symtop EEG system for the Alzheimer’s Disease (AD) data collection.
Parkinson’s Disease (PD) Datasets
Two PD datasets were collected from the University of New Mexico (NMU Center) and the University of Iowa (IU Center), employing a 64-channel Brain Vision EEG system. This system utilized the International 10-10 system for electrode distribution, as depicted in Figure A1, encompassing the following channels: Fp1, Fp2, AF7, AF3, AFz, AF4, AF8, F7, F5, F3, F1, Fz, F2, F4, F6, F8, FT9, FT7, FC5, FC3, FC1, FCz, FC2, FC4, FC6, FT8, FT10, T7, C5, C3, C1, Cz, C2, C4, C6, T8, TP9, TP7, CP5, CP3, CP1, CPz, CP2, CP4, CP6, TP8, TP10, P7, P5, P3, P1, Pz, P2, P4, P6, P8, PO7, PO3, POz, PO4, PO8, O1, Oz, and O2. During EEG data acquisition, the sampling rate was set to 500 Hz. Concurrently, the online reference was configured to channels CPz and Pz for the NMU and IU centers, respectively. Data from these reference channels were excluded from the publicly accessible repository. At the NMU Center, resting-state EEG recordings were obtained from 54 subjects under both eyes-open and eyes-closed conditions. At the IU Center, recordings were conducted with 28 subjects solely under the eyes-open condition. The PD and Normal Control (NC) groups are well-balanced according to the participant numbers across both datasets. To critically evaluate the adaptability of the proposed MACS framework under varying data collection conditions, it was initially trained using data from closed-eye sessions at the NMU center. Subsequently, the performance of the trained model was assessed by applying it directly to open-eye data from the IU center.
Ethical Declaration
The acquisition of EEG data was conducted in strict adherence to ethical guidelines, with approval from the relevant ethical committee, which remains confidential to comply with anonymity rules. Informed consent was duly obtained from all participants prior to their involvement in the study.
A.2. EEG Preprocessing
To ensure efficiency, the raw EEG datasets undergo preprocessing through an automated standard pipeline, as described by Ávila et al. (2023) (Gil Ávila et al., 2023). This pipeline includes filtering, outlier detection, re-referencing, and independent component analysis, designed to eliminate artifacts attributable to low-frequency drifts, high-frequency noise, head movements, cardiac activity, and eye movements. Due to the extensive manual processing and expert monitoring conducted on the AD clinical dataset, we bypassed the standard preprocessing pipeline for this preprocessed data. A total of 63 effective electrodes were employed for the two PD datasets, each utilizing a different reference electrode. For the MCI dataset, 62 effective electrodes were used, while the AD dataset employed 16 electrodes. To ensure temporal resolution consistency, all EEG recordings were uniformly downsampled to 250 Hz. For spatial alignment in the transferability test, linear interpolation was employed before inputting the data into the MACS framework to maintain consistency across the datasets.
Appendix B Implementation Details
B.1. MACS Encoder Architecture
Here, we provide additional details on the MACS Encoder’s two fundamental structures - the convolution starter $g_{str}$ and the temporal clipper $g_{clp}$ - that are omitted from the main body of the text for brevity.
Convolution Starter
The convolution starter $g_{str}$ constitutes a spatial one-dimensional convolution layer and a temporal convolution layer. Each layer is equipped with $d$ kernels, with the spatial layer utilizing kernels of size $(d,1)$ and the temporal layer employing kernels of size $(1,25)$, where $d$ denotes the number of EEG electrodes. We selected a kernel size of 25 sampling points for the EEG’s temporal layer, corresponding to the preprocessed sampling rate of 250 Hz. This kernel size enables the convolutional filter to extract features from a receptive field representing 100 ms of neural activities. The stride for both convolution operations is set to 1, ensuring a comprehensive analysis of the input signals. This architecture is designed to effectively capture generalized spatial and temporal patterns within the EEG data.
Temporal Clipper
The output generated by the convolution starter $g_{str}$ is a feature embedding denoted by $X_{d,T_{s}}$. This embedding is partitioned into $i$ non-overlapping feature clips $\{\ X_{d,t_{i}}\}\ _{i=1}^{I}$ by the temporal clipper $g_{clp}$. Consequently, the chosen value of $I$ plays a critical role in determining the temporal scale at which the dynamic functional network is analyzed in the subsequent manifold-based dynamic attention block $g_{datt}$. For instance, to observe brain synchronization conditions at a one-second temporal scale, $I$ should be configured as $T_{s}/fs$ for the embedding $X_{d,T_{s}}$. Experiments have been conducted to evaluate the effects of the selected hyperparameter, as illustrated in Table 3 of the main text. These experiments ensure that the temporal resolution is appropriately aligned with the desired analytical scale.
B.2. Evaluation Metrics
The metrics used to assess the model’s performance include accuracy, precision, recall, and the F-1 score. Building on the model trained at the fragment level, an ensemble strategy is employed to generate subject-level predictions. This approach involves averaging the probabilities at the fragment level and then determining the subject-level label using the optimal threshold, which is identified by employing Youden’s J statistic  (Youden, 1950). It is crucial to note that this strategy is consistently applied across all comparative methods to ensure fairness in the evaluation process.
Figure A2. Qualitative comparison of the MACS framework with state-of-the-art methods: t-SNE visualization of latent distribution in Parkinson’s disease (PD) data throughout training progress.
Appendix C Extended Experimental Results
C.1. Comparison Study
To guarantee fairness in the comparative analysis, all state-of-the-art (SOTA) methods implemented identical fold divisions, fragment partitioning, cross-validation approaches, and evaluation strategies.
Table A2 delineates the comprehensive results of this comparative study, showcasing full evaluation metrics. While MACS may not achieve the highest scores in every single metric, it demonstrates superior overall performance in the recognition of both MCI and PD, surpassing SOTA methods.
Qualitatively, the MACS framework was compared with three SOTA methods by visualizing the learning progress. This was achieved using t-SNE mappings to depict the embedding features of trusted samples identified by the model across various epochs. Based on their original ground truth labels, before they were reclassified as part of the unreliable annotation sets through label interchanging, the samples identified by the model as trusted were categorized as either appropriately trusted or misplaced. The main text details the model training process on the MCI dataset, showcasing the efficiency of the MACS framework. Complementing this, Figure A2 in the Appendix displays the results obtained from the PD dataset.
For Sel-CL (Li et al., 2022), Promix (Xiao et al., 2023), and our proposed MACS framework, training began at the initial epoch, and thus we visualize the progress starting from epoch 1 with intervals of 5 epochs. For the CTW method (Ma et al., 2023), which includes a warm-up phase before training, t-SNE visualization commenced after the warm-up phases, beginning from epoch 15 with the same 5-epoch intervals.
As the model trains, the desired outcome is that more samples are correctly selected, which enhances representation learning and facilitates the formation of class-specific clusters. The observations on both the MCI dataset (Figure 5 in the main text) and the PD dataset (Figure A2) have demonstrated the efficiency of the MACS framework.
Table A1. Comparative study of state-of-the-art methods for learning with well-annotated and unreliable annotated data.
[MCI] 4-Fold Cross-Validation [PD] 3-Fold Cross-Validation
Parameters Configuration Accuracy Precision Recall F1 Configuration Accuracy Precision Recall F1
0 83.10(6.93) 78.41(17.67) 92.46(8.95) 83.64(9.09) 0 85.18(3.21) 85.46(4.78) 85.19(6.41) 85.15(3.37)
200 87.60(4.46) 87.29(12.65) 86.30(5.96) 86.39(7.58) 50 83.33(0.00) 81.76(5.09) 83.01(5.58) 83.30(0.93)
300 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 100 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
400 86.31(5.55) 85.36(12.25) 85.60(13.86) 84.71(9.69) 200 85.18(3.21) 89.17(10.10) 81.48(6.41) 84.69(2.61)
Memory Length 500 85.33(4.52) 84.67(13.37) 84.21(4.80) 84.07(7.77) 300 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
2s 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 1s 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
1s 87.65(4.31) 87.12(10.98) 87.26(10.90) 86.75(7.87) 500ms 81.48(3.20) 89.17(10.10) 74.08(16.97) 79.33(6.90)
Temporal Scale 500ms 74.06(6.09) 67.24(13.56) 94.36(7.86) 77.60(8.15) 250ms 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
15 86.46(3.87) 93.03(8.13) 76.92(18.90) 82.69(11.65) 8 83.33(0.00) 85.00(4.33) 81.48(6.41) 82.97(1.07)
20 87.60(4.46) 82.68(13.28) 93.94(7.01) 87.39(7.30) 12 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
25 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 16 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
K Value 30 87.60(4.46) 82.96(12.50) 93.94(7.01) 87.61(6.87) 20 85.18(3.21) 93.33(11.55) 77.78(11.11) 83.90(3.76)
Table A2. The class-discriminative activation heatmaps generated by the MACS framework highlight localized Regions of Interest (ROIs) for detecting abnormalities in EEG signals under MCI, AD, and PD conditions. These heatmaps were averaged across all samples within each dataset, including the MCI, AD, and PD (from both the NMU and IU centers) datasets. The scatter plots illustrate the consistency between each sample and the averaged results, statistically quantified by the Spearman coefficient ($^{*}/CC^{**}$, indicating meaningful correlation) and the corresponding p-values ($^{*}/pp^{**}$, signifying significance).
[MCI] 4-Fold Cross-Validation [PD] 3-Fold Cross-Validation
Parameters Configuration Accuracy Precision Recall F1 Configuration Accuracy Precision Recall F1
0 83.10(6.93) 78.41(17.67) 92.46(8.95) 83.64(9.09) 0 85.18(3.21) 85.46(4.78) 85.19(6.41) 85.15(3.37)
200 87.60(4.46) 87.29(12.65) 86.30(5.96) 86.39(7.58) 50 83.33(0.00) 81.76(5.09) 83.01(5.58) 83.30(0.93)
300 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 100 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
400 86.31(5.55) 85.36(12.25) 85.60(13.86) 84.71(9.69) 200 85.18(3.21) 89.17(10.10) 81.48(6.41) 84.69(2.61)
Memory Length 500 85.33(4.52) 84.67(13.37) 84.21(4.80) 84.07(7.77) 300 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
2s 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 1s 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
1s 87.65(4.31) 87.12(10.98) 87.26(10.90) 86.75(7.87) 500ms 81.48(3.20) 89.17(10.10) 74.08(16.97) 79.33(6.90)
Temporal Scale 500ms 74.06(6.09) 67.24(13.56) 94.36(7.86) 77.60(8.15) 250ms 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
15 86.46(3.87) 93.03(8.13) 76.92(18.90) 82.69(11.65) 8 83.33(0.00) 85.00(4.33) 81.48(6.41) 82.97(1.07)
20 87.60(4.46) 82.68(13.28) 93.94(7.01) 87.39(7.30) 12 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
25 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 16 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
K Value 30 87.60(4.46) 82.96(12.50) 93.94(7.01) 87.61(6.87) 20 85.18(3.21) 93.33(11.55) 77.78(11.11) 83.90(3.76)
Figure A3. The class-discriminative activation heatmaps generated by the MACS framework highlight localized Regions of Interest (ROIs) for detecting abnormalities in EEG signals under MCI, AD, and PD conditions. These heatmaps were averaged across all samples within each dataset, including the MCI, AD, and PD (from both the NMU and IU centers) datasets. The scatter plots illustrate the consistency between each sample and the averaged results, statistically quantified by the Spearman coefficient ($^{*}/CC^{**}$, indicating meaningful correlation) and the corresponding p-values ($^{*}/pp^{**}$, signifying significance).
C.2. Hyperparameter Tuning Study
Memory Length
The multi-view contrastive loss is a pivotal component in our MACS framework, as it significantly depends on the adequacy of positive and negative pairs for effective contrastive learning. However, the prevalent scarcity of clinical data poses a substantial challenge. To circumvent the constraints imposed by batch size and to capitalize on the more extensively available data during the learning process, we adopt a storage strategy as outlined by (Ortego et al., 2021). This strategy involves retaining the preceding $M$ features in memory, with $M$ denoting the memory length. As illustrated in Table A2, the influence of memory length on the model’s performance was assessed for both MCI and PD datasets via an N-fold cross-validation experiment. This hyperparameter tuning study determined the optimal memory lengths to be 300 for the MCI dataset and 100 for the PD dataset. The results reveal a bell-shaped relationship between memory length and model performance, suggesting an optimal point appears to correlate with the dataset size, as evidenced by the MCI dataset containing approximately three times more EEG fragments than the PD dataset.
k-Nearest Neighbors
The parameter $K$ determines the number of nearest neighbors selected to infer potential labels, using the k-Nearest Neighbors algorithm in the Stratifier $f_{St}$ module of the MACS framework. A thorough exploration of $K$, employing N-fold cross-validation, is summarized in Table A2. In the experiments, the minimum value of $K$ and the interval are contingent upon the batch size of MCI and PD datasets. Similar to memory length, $K$ exhibits an ’inverted-U relationship’. Thus, $K$’s optimal configuration should be well-coordinated with the batch sizes.
Temporal Scale
As delineated in Appendix B.1, the variable $I$ dictates the temporal scale employed for analyzing characteristics within the dynamic functional networks. To this end, a series of experiments were conducted, exploring the effects of temporal scales, extending from the millisecond level to the second level. The findings presented in Table A2 indicate that adopting a second-level temporal scale is necessary for achieving robust and improved outcomes. Specifically, the temporal scale was configured to 2 seconds for the MCI dataset and 1 second for the PD dataset. The decision to configure the maximum temporal scale for PD data at 1 second was informed by the minimal length of PD data, which stands at 2 seconds. This configuration was strategically chosen to ensure that a minimum of one sample is capable of yielding two distinct data segments, thus meeting the structural prerequisites of our model. The emergence of more effective EEG markers at the larger temporal scales, as presented in Table A2, may be attributable to the ’slowing’ phenomenon in brain activity, which is more prominently manifested at broader temporal scales in patients with neurodegenerative diseases (D’Atri et al., 2021; Jones et al., 2022)
C.3. Region of Interest Localized by MACS
We applied a class-discriminative activation algorithm inspired by the Grad-CAM method (Selvaraju et al., 2017) to localize the region of interest (ROI) identified by the MACS framework while analyzing abnormalities in MCI, AD, and PD based on EEG signals.
To enhance our understanding, we computed the average activation across all samples within each dataset, resulting in the generation of four ROI heatmaps corresponding to MCI, AD, PD (MNU center), and PD (IU center). Concurrently, we aimed to assess the uniformity of the sample-specific responses relative to their respective averaged ROI activation maps. To achieve this, we utilized the Spearman rank correlation method, as detailed in the referenced pipeline  (Song et al., 2022), calculating the correlation coefficients between the averaged activation maps and each sample, along with an overall quantification metric termed $ROI_{f}$. The higher $ROI_{f}$ indicates greater consistency between all samples and their average result.
Here, a Spearman correlation coefficient ($C_{Spearman}$) of 1 (or -1) indicates a perfect positive (or negative) correlation, while a coefficient of 0 signifies the absence of correlation. In each heatmap’s right panel, the visualization includes blue triangle markers that identify individual samples. The correlation significance within these heatmaps is demarcated by two lines: a dotted green line at $C_{Spearman}=0.3$ differentiates between negligible and meaningful correlations, and a solid green line at $C_{Spearman}=0.5$ highlights correlations ranging from moderate to perfect. To establish the statistical relevance of these correlations, p-values were calculated. Correlations with p-values below 0.05 are marked by a solid red line, indicating statistical significance, while a dotted red line at p = 0.01 delineates a more rigorous threshold for significance.
Similar patterns are observed in conditions of MCI and AD, notably with significant ROIs in the parietal-occipital and frontal-temporal lobes. This may be indicative of a pathological continuum from MCI to AD, with MCI being an earlier stage in the progression towards the more pronounced dementia typical of AD. This evidence supports that the MACS framework effectively captures the core patterns in AD, potentially providing EEG evidence of compensatory mechanisms in the early stage (Gaubert et al., 2019). Analogous ROIs, such as the right frontal-temporal lobes and left temporal-parietal lobes, are observed in PD conditions, despite the data originating from two different centers. This observation suggests that the model is capable of capturing shared features across diverse datasets, thereby enhancing its transferability.
C.4. Speed and Parallelism
We assessed the latency of our framework compared to state-of-the-art (SOTA) frameworks, as shown in Table A3. Although our framework is conceptualized as a dual-branch system, the parameters are shared between modules, as illustrated in Figure 2. This dual-branch concept is realized through our contrastive losses, enabling parallelism. To ensure fair speed comparisons in weakly supervised learning, we substituted the encoders in SOTA frameworks with ours. MACS sacrifices approximately 8 ms compared to CTW and Promix to enhance performance robustness through neighbor-based confidence evaluation in the Stratifier. We doubled the speed of MACS compared to Sel-CL by implementing a selection mechanism in the Switcher that reduces pairwise samples for contrastive learning. Additionally, we compared our Encoder with the original encoders in the SOTA frameworks to demonstrate its effectiveness in reducing computational costs. Overall, our framework achieves millisecond-level inference speeds, ensuring practical application and deployment.
Table A3. Comparison of Framework Latency, Encoder Size, and Encoder Computing Complexity.
Frameworks CTW Promix Sel-CL MACS
Latency (ms) 7.39(1.42) 8.61(0.67) 37.35(1.89) 16.2(1.22)
Encoders FCN Dual ResNet ResNet Manifold Attention
Size (MB) 1.01 4.55 2.77 0.95
FLOPs (G) 0.13 0.27 1.10 0.05
Appendix D Data Availability
The PD datasets are publicly accessible and can be downloaded from http://predict.cs.unm.edu/. Access to the MCI and AD datasets will be granted upon request to the corresponding authors and subject to the approval of the collaborating hospitals.
Appendix E Code Availability
To facilitate the reproduction of our experiment, we have made the core code available at https://github.com/ICI-BCI/EEG-MACS.